OpenStudy (anonymous):

use the distributive property to rewrite the expression. (x+3)5 simplify the answer

OpenStudy (anonymous):

u kn dis..... law is like a(b+c) = ab+ac

OpenStudy (zzr0ck3r):

its the same as 5(x+3)

OpenStudy (zzr0ck3r):

a(b+c) = ab+ac so 5(x+3) = 5x + (5*3) = 5x+15